• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Camsonne

    Log entry time 08:15:01 on April 9,2007

    Entry number 198351

    keyword=Owl shift summary

    Shift summary
    00:00 Shift starts run 2231
    2:08 Moving to dummy target
    2:14 Dummy Target in place

    2:50 Noticed CODA dead time of 100 %
    3:00 Start of run 2240 should be a good run

    4:30 End of dummy run
    Alarm in the vaccum of the scattering chamber pressure went up to 3e-5 torr. It seems to be correlated with a small glitch in the cryo.

    4:44 He3 target in place
    asked for 40 uA no change in target variables
    went back to 83 uA,

    ~5:20 another alarm for vaccum Hall A 4K flows seems to reach 25 g/s when it happens while it usually is around 30 g/s, the CHL4K is stable around 20.5 g/s

    6:50 Coda crashed

    7:40 ROC2 crashed needed access to power cycle the crate

    8:00 pass over to next shift

    Run Events
    2231 1.4M He3
    2232 1.4M He3
    2233 1.4M He3
    2234 1.4M He3
    2235 1.4M He3
    2236 1.4M He3
    2237 1.4M He3
    ----------------

    2238 Dummy Coda 100% dead
    2239 Dummy Junk
    2240 Dummy 1.4 M
    2241 Dummy 1.4 M
    2242 Dummy 1.4 M
    2243 Dummy 1.4 M

    -------------------
    2244 1.4M He3
    2245 1.4M He3
    2246 1.4M He3
    2247 1.4M He3
    2248 1.4M He3
    2249 1.4M He3
    2250 1.4M He3 Coda crashed
    2251 1.4M He3
    2252 ROC2 crashed
    2253 test
    2254